# Bridging the Social & Technical Divide in Augmentative and Alternative Communication (AAC) Applications for Autistic Adults

### v1

@@ Abstract.
Natural Language Processing (NLP) techniques are being used more frequently to improve high-tech Augmentative and Alternative Communication (AAC), but many of these techniques are integrated without the inclusion of the users’ perspectives. As many of these tools are created with children in mind, autistic adults are often neglected in the design of AAC tools to begin with. We conducted in-depth interviews with 12 autistic adults to find the pain points of current AAC and determine what general technological advances they would find helpful. We found that in addition to technological issues, there are many societal issues as well. We found 9 different categories of themes from our interviews: input options, output options, selecting or adapting AAC for a good fit, when to start or swap AAC, benefits (of use), access (to AAC), stumbling blocks for continued use, social concerns, and lack of control. In this paper, we go through these nine categories in depth and then suggest possible guidelines for the NLP community, AAC application makers, and policy makers to improve AAC use for autistic adults.
accessibility, augmentative and alternative communication, AAC, ASD, autism, NLP, word prediction, large language models, LLMs, ChatGPT
copyright: acmlicensedjournalyear: 2024ccs: Human-centered computing Accessibility technologiesccs: Human-centered computing Accessibility systems and toolsccs: Human-centered computing Accessibility design and evaluation methodsccs: Human-centered computing Natural language interfacesccs: Human-centered computing Empirical studies in HCIccs: Social and professional topics People with disabilities
@@ 1. Introduction
Thanks to the neurodiversity movement online (Kapp, 2020; Botha et al., 2024), many people who were previously overlooked for diagnoses as children are now realizing that they are autistic as adults. Since these autistic individuals are coming into their diagnoses well into adulthood, they are learning about accommodation much later in life. This means that tools like Augmentative and Alternative Communication (or AAC), which replace or supplement a person’s oral communication, are discovered by people who 1) already know how to read and write, 2) have been speaking orally (to their detriment) for years (Zisk and Dalton, 2019), and 3) may not have access to professionals such as speech-language pathologists. Furthermore, the tools themselves are usually developed with children in mind, with speech-language pathologists teaching children how to use AAC to speak their first language.
@@
Even though there has been recent recognition within AAC research showing that AAC use can build agency (Valencia et al., 2020) and garner empathy from the people in the AAC user’s life (van Grunsven and Roeser, 2022), autistic adults are largely ignored within AAC research (Ganz et al., 2012; Lorah et al., 2022; Trembath et al., 2014) and very few are involved in the research of new AAC tools (Bastable et al., 2021). Because of this, autistic adults have learned to navigate a world of AAC tools that are not made for them.
@@
On the technological side of things, research in high-tech AAC tools [^1] AAC tools can range across the spectrum of technology levels—from writing, spelling out words on a letter-board, or using cards with pictures on them; to tablet/phone-based applications that cost several hundreds of dollars. has seen a recent uptick in interest from the Natural Language Processing (NLP) community due to the introduction of large language models (LLMs) like ChatGPT to the general public (Cai et al., 2022; Kumar et al., 2022; Valencia et al., 2023; Yusufali et al., 2023). However, we should tread carefully as NLP and other machine learning techniques can cause harm if not implemented correctly. There are crucial privacy issues such as training on users’ data with out their knowing or sending/saving data to “the cloud” with questionable levels of security. Furthermore, implementing NLP techniques blindly without the inclusion of the community of users can backfire. Thankfully, there has been a push to include AAC users as contributors in AAC development (Bastable et al., 2021; Beneteau, 2020). Our work continues down this path by beginning conversations with AAC users to bridge the gap between what can be done technologically and what users actually want, with a focus on high-tech, computer-based AAC applications.
@@
In this paper, we wanted to learn about the needs of autistic adults such that we can improve AAC applications and dig into some of the deeper issues as to why AAC tools are not always available. In the vein of looking to improve AAC tools for a sub-community of users who are often unheard—autistic adults, we were interested in the following research questions:
@@
  • RQ1: What are pitfalls and challenges that autistic AAC users run into when using existing AAC applications?
@@
  • RQ3: How does the person’s identity as an autistic person factor into their AAC use?
  • RQ4: How do participants feel about NLP techniques being integrated into their AAC apps?
To learn more about this intersection of identities, we ran 12 in-depth semi-structured interviews with autistic adults who were recruited online. We outline some themes that we have identified throughout the interviews and provide design recommendations to address them from the perspective of three different societal levels. To match the language autistic adults use (Zisk and Konyn, 2022), we will be using the terms ”speaking orally”, ”using mouth-words”, or ”speaking” (vs ”nonspeaking”) interchangeably throughout the paper to refer to using one’s vocal tract to produce language.
2. Related Work
@@ 2.1. Autistic Adult AAC Users
As mentioned in the Introduction, the majority of AAC work either involves autistic children or adults with physical disabilities. However, there is a recent (although small) cohort of work that has recognized this intersectional identify of adult autistic users of AAC (Donaldson et al., 2021; Hajjar and McCarthy, 2022; Holyfield et al., 2020; Hughes et al., 2022; Zisk and Dalton, 2019). Even so, a lot of this work has a limited context. For example, Holyfield et al. (2017) performed a literature review of introductions of AAC (“interventions”) into the lives of autistic adolescents and adults. Out of the 19 participants mentioned in the papers they found, only 4 of them were adults.
Other works have focused on adults with intellectual and developmental disabilities broadly (Holyfield et al., 2020; Hughes et al., 2022), which includes people with autism spectrum disorder. Previous work includes using AAC to improve users’ skills such as single word sight reading (Holyfield et al., 2020) or achieving their own personalized communication goals (Hughes et al., 2022). Hajjar and McCarthy (2022) asked adult AAC users how they experienced using AAC during physical activities like sports.
@@
To our knowledge, there have only been two papers written so far looking specifically at autistic adults’ general AAC use (Zisk and Dalton, 2019; Donaldson et al., 2021)—both papers focused on speaking adults. Our study mirrors these works since we looked at autistic adults who were not “completely nonspeaking”. (However, we speculate that it is probably difficult for someone to consider themselves fully nonspeaking if they have forced speech for most of their life and only encountered AAC as an adult, as many of our participants have.) Zisk and Dalton (2019) reviewed academic and non-academic sources to find recurring themes of how speaking autistic adult AAC users communicate when they can occasionally use speech, what types of tools they use to communicate (both specialized for AAC and not), and any barriers to using AAC. Donaldson et al. (2021) performed an open-ended written survey with six participants, asking questions specific to AAC use history, access, attitudes to use, and any recommendations they might have for people who might use or interact with people who use AAC. While our work has rediscovered some of the same social themes (e.g., choice of AAC, autonomy, acceptance), we also dive into technical and privacy themes.
2.2. Technological Advances in AAC
Technologically speaking, most research on improving high-tech AAC tools focuses on customizablity or automating/predicting what the user means to do (Curtis et al., 2022). Much of the prediction involves predicting the text that the user would want to say. In the field of natural language processing (NLP), this primarily involves the use of a “language model”. Language modeling is a way of creating a probabilistic model of text. That is, after training on many documents, they can learn to predict what word(s) will come next given the words it has seen so far. For AAC, research has focused on next word prediction (Adhikary et al., 2019; Freedman et al., 2013; Schadle, 2004; Trnka et al., 2007; Wiegand and Patel, 2012), contextual or discourse-based prediction (Freedman et al., 2013; Mitchell and Sproat, 2012; Perkoff, 2021), personalized language (Di Mascio et al., 2019; Freedman et al., 2013; Li et al., 2022; Valencia et al., 2023), or generating sentences from a set of keywords (Cai et al., 2022; Kumar et al., 2022). Other unique approaches include using automatic speech recognition (speech-to-text) to improve word prediction by listening to AAC users’ conversation partners and using the recognized words as additional context (Adhikary et al., 2019); and classifying emotions based on self-labeled non-lexical vocalizations (Johnson et al., 2023).
Using earlier language modeling techniques, researchers would get their data from a variety of sources, including children’s stories (Mühlenbock and Lundälv, 2011), crowdsourced conversations (Vertanen and Ola Kristensson, 2011; Vertanen, 2017), online sources like Reddit (Yusufali et al., 2023), and TV shows (Yusufali et al., 2023). Recently, large language models (LLMs), such as ChatGPT, have been used for speeding up the writing process for AAC (Cai et al., 2022; Kumar et al., 2022; Valencia et al., 2023; Yusufali et al., 2023). AAC users are enthusiastic but cautious about employing LLMs to assist with their communication (Valencia et al., 2023). LLMs are extremely capable of generating coherent text, but there can still be issues of controlling the model to generate exactly what you want, especially if you are not used to using LLMs. Despite these advances, very little work has been done on non-English AAC (Amery et al., 2022; Mühlenbock and Lundälv, 2011).
There are also recent improvements in other aspects of AAC beyond the language component. The research in these areas include finding better interfaces (Gevarter et al., 2020; Light et al., 2019; Yusufali et al., 2023), improving the robustness of apps on smart devices (Laxmidas et al., 2021; Roark and Gutkin, 2022), and incorporating additional modalities such as physiological sensors (Holyfield and O, 2022), gestures (Ascari et al., 2021), or photographs (Fontana de Vargas and Moffatt, 2021; Pitt and McCarthy, 2021) and pictograms (Martínez-Santiago et al., 2016; Sampath et al., 2012) in addition to or instead of text.
@@ 3. Methods
    Questions asking the participant to recall a recent conversation where they used AAC and describe details about how that conversation went.
@@
    Questions asking the user to think of particular scenarios when they used AAC.
@@
    Questions for describing a typical day of using AAC and what types of tools they used.
@@
    Questions asking about when they have felt empower or disempowered because of their AAC.
@@
    Questions about how they select what AAC to try.
@@
    Questions about privacy (ending with a privacy table).
The privacy table was a list of potential features that were presented to the participant after they had a chance to think about the privacy of their AAC on their own. For each of the seven features we came up with, the participant had to rate whether they would allow it in an AAC application they would use. More information about the table and its results are presented in Section 5.9.
We recruited participants via the Facebook group “Ask Me, I’m an AAC user!” and Twitter. The participants were asked to fill out an initial intake questionnaire (Appendix Section A), and if they qualified, they were emailed the consent form and a video of the first author reading the consent form aloud [^2] This practice of recording the consent form was taken with permission from Charlotte Brooks, which was well-received by autistic adult participants.. If they consented, they were asked to schedule an interview and were given a PDF containing all the interview questions & a PDF with the privacy table—the latter of which they were told not to open until the appropriate time. Participants were allowed to script their answers to the questions ahead of time, and were especially encouraged to do so if they had planned to use AAC during the interview.
We kept the recruiting open until we had 12 participants  [^3] It has been shown (Ando et al., 2014) that themes begin to stabilize after around 12 participants.. Seventy-eight people signed up for the study and filled out the initial intake questionnaire, twenty-five people were invited to participate in the interview, and twelve people were interviewed. For confidentiality, we have carefully revised the anecdotes to eliminate any information that could potentially identify the participants.
@@
Most of the interviews were performed via Zoom with video and audio. The interviews lasted about an hour, with the average being 1 hour and 18 minutes (standard deviation of 31.2 minutes, minimum 50 minutes, maximum 2 hours & 40 minutes). Interviewees were encouraged to use their AAC during the interview. One interview was conducted entirely asynchronously via email & text document. Another interview was conducted half on Zoom (cut off at 1 hour) and half asynchronously. For consistency, all interviews were run by the first author, who was careful to note the language the participant used when talking about how they referred to the use of mouth-words.
@@ 3.0.1. Analysis
We analyzed the interviews through thematic analysis (Braun and Clarke, 2006). We used Zoom’s built-in automated transcription tool to get an initial draft of the transcript. The transcripts were then corrected and verified by the authors. If an important section of the interview could not be confidently transcribed, the authors asked the interviewee for clarification via email later on. The authors independently formulated an initial set of themes by analyzing the transcripts in an asynchronous manner. Subsequently, they engaged in discussions to merge their respective themes, followed by a process of refining these combined themes and evaluating their alignment with the transcript content. Through iterative revisions and affinity diagramming, the authors reached a consensus on the final set of categories and themes presented in this paper.
4. Types of AAC Applications
There appears to be a divide between two categories of AAC applications based on input type: those that use typing-based input and those that use symbol-based input. Typing-based apps enable users to input any word they want on the fly, while symbol-based apps require pre-programming of what words will be displayed on buttons and what the layout of the buttons will be. The divide seems to revolve around two commonly-used applications (Proloquo2Go and Proloquo4Text), both developed by the company AssistiveWare [^4] https://www.assistiveware.com/. During the course of this research, AssistiveWare has come out with a new application in the Proloquo suite simply called Proloquo. Proloquo seems to solve some of the problems of this divide by beginning to merge the functionality of Proloquo2Go and Proloquo4Text. Proloquo is now a button-based layout similar to Proloquo2Go but without symbols (only words) and with more predictive text options. Proloquo4Text consists primarily of on-screen typing with the option to save phrases.

### v3

# Aging Up AAC: An Introspection on Augmentative and Alternative Communication Applications for Autistic Adults

@@ Abstract.
High-tech Augmentative and Alternative Communication (AAC) has been rapidly advancing in recent years due to the increased use of large language models (LLMs) like ChatGPT, but many of these techniques are integrated without the inclusion of the users’ perspectives. Autistic adults have been particularly neglected in the design of AAC tools. We conducted in-depth interviews with 12 autistic adults to find the pain points of current AAC and determine what technological advances they might find helpful. We found 8 different categories of themes from our interviews: input flexibility, output flexibility, selecting or adapting AAC, contexts for AAC use, benefits, access as an adult, stumbling blocks for continued use, and control of communication. In this paper, we go through these categories in depth – comparing each to prior work – and then highlight novel findings to suggest possible research directions.
accessibility, augmentative and alternative communication, AAC, ASD, autism, autistic adults, LLMs, word prediction
conference: Preprint; 2025; arXivccs: Human-centered computing Accessibility technologiesccs: Human-centered computing Accessibility systems and toolsccs: Human-centered computing Accessibility design and evaluation methodsccs: Human-centered computing Natural language interfacesccs: Human-centered computing Empirical studies in HCIccs: Social and professional topics People with disabilities
Figure 1 is a collage of various AAC applications used by participants. In the collage there is: 1. A picture of a hand holding a phone that shows the app My Voice. 2. A screenshot of someone’s homeboard in CoughDrop. 3. A picture of someone typing on a keyboard in front of a monitor with a tablet in their lap. 4. A screenshot of someone typing a sentence in Proloquo. 5. A picture of a tablet showing TD Snap.
Figure 1. A compilation of our participants’ various AAC devices and interfaces. The applications (from top to bottom, left to right) are My Voice, CoughDrop, ClaroCom (with keyboard setup), Proloquo, and TD Snap.
@@ 1. Introduction
Thanks to the neurodiversity movement online (Kapp, 2020; Botha et al., 2024), many people who were previously overlooked for diagnoses as children are now realizing that they are autistic as adults. Since these autistic individuals come into their diagnoses well into adulthood, they are learning about accommodation much later in life (Donaldson et al., 2021). This means that tools like Augmentative and Alternative Communication (or AAC), which replace or supplement a person’s oral communication, are discovered by people who 1) already know how to read and write, 2) have been speaking orally (to their detriment) for years (Zisk and Dalton, 2019), and 3) may not have access to professionals such as speech-language pathologists. Furthermore, the tools themselves are usually developed with children in mind, with speech-language pathologists teaching children how to use AAC to speak their first language.
@@
Although there has been recent recognition within AAC research showing that AAC use can build agency (Valencia et al., 2020) and garner empathy from people in the AAC user’s life (van Grunsven and Roeser, 2022), autistic adults are largely ignored within AAC research (Ganz et al., 2012; Trembath et al., 2014; Lorah et al., 2022), especially when designing new AAC tools (Bastable et al., 2021). Because of this, autistic adults have learned to navigate a world of AAC tools that are not made for them.
@@
Technologically-speaking, research on high-tech AAC tools [^1] AAC tools can range across the spectrum of technology levels—from writing, spelling out words on a letter-board, or using cards with pictures on them; to tablet/phone-based applications that cost several hundreds of dollars. has seen a recent uptick in interest from the Natural Language Processing (NLP) community due to the introduction of large language models (LLMs) like ChatGPT to the general public (Cai et al., 2022; Kumar et al., 2022; Valencia et al., 2023; Yusufali et al., 2023a). However, we should tread carefully as NLP and other machine learning techniques can cause harm if not implemented correctly (Huang et al., 2023). There are crucial privacy issues such as training on users’ data without their knowledge (Crawford et al., 2024) or sending/saving data to “the cloud” with questionable levels of security (Hamidi et al., 2018). Furthermore, implementing NLP techniques blindly without the inclusion of the community of users can backfire. Thankfully, there has been a push to include AAC users as contributors in AAC development (Kudryashov, 2021; Bastable et al., 2021; Beneteau, 2020; Taylor and Balandin, 2020). Our work continues down this path by beginning conversations with AAC users to bridge the gap between what can be done technologically and what users actually want, with a focus on high-tech, computer-based AAC applications.
@@
In this paper, we characterize the needs of autistic adults in order to improve AAC applications and investigate deeper issues as to why some adults do not use AAC applications. In the vein of looking to improve AAC tools for a sub-community of users who are often unheard—autistic adults, we were interested in the following research questions:
@@
  • RQ1: What are the benefits and challenges that autistic AAC users face when using existing AAC applications?
@@
  • RQ3: How does a person’s identity as an autistic person factor into their AAC use?
  • RQ4: How do autistic adults feel about various technical considerations for future AAC applications (i.e., the integration of NLP techniques)?
To learn more about this intersection of identities, we conducted 12 semi-structured in-depth interviews with autistic adults who were recruited online. We outline themes that we have identified throughout the interviews and compare our findings to prior work  [^2] Many social issues regarding AAC use arose in addition to technological issues. We will not be discussing the social issues since they are outside the scope of this paper. If this is relevant to your work and you would like to know what we found, please contact the first author.. Finally, we highlight potential new research directions based on our interviews.
To match the language autistic adults use (Zisk and Konyn, 2022), we will be using the terms ”speaking orally”, ”using mouth-words”, or ”speaking” (vs. ”nonspeaking”) interchangeably throughout the paper to refer to using one’s vocal tract to produce language.
2. Previously-Identified Issues with AAC
@@ 2.1. Autistic Adult AAC Users
Some work with intellectual and developmental disabilities broadly, including people with autism spectrum disorder, has looked at how AAC can improve users’ skills such as single word sight reading (Holyfield et al., 2020) or achieving their own personalized communication goals (Hughes et al., 2022). These types of work track how well participants use AAC rather than how they feel about the technology. To our knowledge, there have only been two papers written so far looking specifically at autistic adults’ general AAC use (Zisk and Dalton, 2019; Donaldson et al., 2021)—both papers focused on speaking adults. Our study mirrors these works since we interviewed autistic adults who were not “completely nonspeaking”. (We speculate that it is difficult for someone to consider themselves fully nonspeaking if they have forced speech for most of their life and only encountered AAC as an adult, as many of our participants have.)
Zisk and Dalton (2019) reviewed academic and non-academic sources to find recurring themes of how speaking autistic adult AAC users communicate when they can occasionally use speech, what types of tools they use to communicate (both specialized for AAC and not), and any barriers to using AAC. They mention that autistic adults are unaware of AAC options or ways of getting support, and that lacking an autism diagnosis can impede adults from getting the aid they need. Zisk and Dalton (2019) also identify the cost of applications or preconceived notions of AAC as issues for autistic adults to acquire AAC. Donaldson et al. (2021) performed an open-ended written survey with six participants, asking questions specific to AAC history, access, and attitudes toward AAC use. They also asked participants for any recommendations for people who might use AAC or interact with people who use AAC. While our work has rediscovered some of the same social themes as Donaldson et al. (2021) (e.g., choice of AAC, autonomy), we also dive into technical and privacy themes. Donaldson et al. (2021) also focus more on discovering AAC as an adult, acceptance of AAC use, and forcing speech. These papers will be referenced throughout this work.
2.2. Other Populations of AAC Users
There are some fairly universal issues with AAC use among adult users. Some recurring themes across the literature include issues with the speed of communication and conversation pacing (Higginbotham et al., 2016; Kane et al., 2017; Waller, 2019; Shin et al., 2020), access to AAC (Waller, 2019; Dai et al., 2022; Amery et al., 2022; Pope et al., 2022; Donaldson et al., 2023), being sufficiently trained in using the AAC (Waller, 2019; Gibson et al., 2020), low-quality speech synthesis (Beukelman et al., 2007; Kane et al., 2017; Gibson et al., 2020), and difficulties finding words (Shin et al., 2020; Obiorah et al., 2021). Past participants mentioned a tendency for AAC devices to be too focused on needs rather than other topics of conversation (Kane et al., 2017; Waller, 2019; Obiorah et al., 2021; Dai et al., 2022) or lack personality/humor (Kane et al., 2017; Dai et al., 2022; Weinberg et al., 2024). There are also impacts on the agency (both positive and negative) of the AAC user (Ibrahim et al., 2018; Valencia et al., 2020; Hajjar and McCarthy, 2022; Donaldson et al., 2023). Compared to these past works that focus primarily on the pain points of existing technologies, we ask users to reflect on issues of existing AAC technologies while also considering how future AAC might be made. In particular, we orient participants to think about empowerment, NLP/AI use, and privacy.
There have also been studies including specific populations of non-autistic (i.e., allistic) adults who use AAC. In addition to issues mentioned in the previous paragraph, Gibson et al. (2020) found adults with mild intellectual disabilities want customizable interfaces, tutorials on how to use the app, fewer options presented at a time, images to explain button functionality, and audio playback of text. (Obiorah et al., 2021) saw that adults with aphasia would like to integrate other digital devices into their AAC use. Kane et al. (2017) worked with adults with ALS, who experienced difficulties with volume and switching roles in a conversation.
Very little work directly involves children with disabilities in the design of assistive technologies (Benton and Johnson, 2015; Spiel et al., 2019). Instead, much of the work has a clinical focus—looking at how effective the child’s communication is (Logan et al., 2017), or asking parents/guardians, teachers, etc. of the children to describe the child’s AAC use (Beneteau, 2020). That said, Shin et al. (2020) have noted that autistic children have trouble generating full sentences with their AAC and their parents have difficulty modeling sentence creation with symbols. Walker and Chung (2022) saw that autistic children more often use AAC with adults than with other children. Most autistic children have a preference for high-tech AAC (vs low-tech AAC or signing) (Lorah et al., 2022). Throughout this paper, we also mention previous work relevant to the corresponding themes.
@@ 3. Methods
    Questions asking the participant to recall a recent conversation where they used AAC and describe details about how that conversation went (RQ1).
@@
    Questions asking the user to think of particular scenarios when they used AAC (RQ1).
@@
    Questions for describing a typical day of using AAC and what types of tools they used (RQ1,RQ2).
@@
    Questions asking about when they have felt empowered or disempowered because of their AAC (RQ3).
@@
    Questions about how they select what AAC to try (RQ2).
@@
    Questions about privacy (ending with a privacy table) (RQ4).
The privacy table was a list of potential NLP features that were presented to the participant after thinking about the privacy of using AAC on their own. For each of the seven features we came up with, the participant had to rate whether they would allow it in an AAC application they would use. More information about the table and its results is presented in Section 5.8. The research design, as well as the consent form, were approved by the IRB board of the University of Pennsylvania.
We recruited participants via the Facebook group “Ask Me, I’m an AAC user!” and X (formerly Twitter). The participants were asked to fill out an initial intake questionnaire (Appendix Section A), and if they qualified, they were emailed the consent form and a video of the first author reading the consent form aloud [^3] This practice of recording the consent form was taken with permission from Charlotte Brooks, which was well-received by autistic adult participants.. If they consented, they were asked to schedule an interview and were given a PDF containing all the interview questions & a PDF with the privacy table—the latter of which they were told not to open until the appropriate time. Participants were allowed to script their answers to the questions ahead of time, and were especially encouraged to do so if they had planned to use AAC during the interview.
We kept the recruiting open until we had 12 participants  [^4] It has been shown (Ando et al., 2014) that themes begin to stabilize after around 12 participants.. Seventy-eight people signed up for the study and filled out the initial intake questionnaire, twenty-five people were invited to participate in the interview, and twelve people were interviewed. For confidentiality, we have revised the anecdotes to eliminate any information that could potentially identify the participants.
@@
Most of the interviews were performed via Zoom with video and audio. The interviews lasted about an hour, with the average being 1 hour and 18 minutes (standard deviation of 31.2 minutes, minimum 50 minutes, maximum 2 hours & 40 minutes). Interviewees were encouraged to use their AAC during the interview. One interview was conducted entirely asynchronously via email & text document. Another interview was conducted half on Zoom (cut off at 1 hour) and half asynchronously. For consistency, all interviews were run by the first author.
@@ 3.0.1. Analysis
We analyzed the interviews through thematic analysis (Braun and Clarke, 2006). We used Zoom’s built-in automated transcription tool to get an initial draft of the transcript. The transcripts were then corrected and verified by the authors. If an important section of the interview could not be confidently transcribed, the authors asked the interviewee for clarification via email. The authors independently formulated an initial set of themes by analyzing the transcripts in an asynchronous manner. Subsequently, they engaged in discussions to merge their respective themes, followed by a process of refining these combined themes and evaluating their alignment with the transcript content. Through iterative revisions and affinity diagramming, the authors reached a consensus on the final set of categories and themes presented in this paper.
4. Categories of High-Tech AAC Applications
We observed a divide between two categories of AAC applications based on input type: those that use typing-based input and those that use symbol-based input. Typing-based apps enable users to input any word they want on the fly, while symbol-based apps require pre-programming of what words will be displayed on buttons and what the layout of the buttons will be. The divide seems to revolve around two commonly-used applications (Proloquo2Go and Proloquo4Text), both developed by the company AssistiveWare [^5] https://www.assistiveware.com/. During the course of this research, AssistiveWare had come out with a new application in the Proloquo suite simply called Proloquo. Proloquo seems to solve some of the problems of this divide by beginning to merge the functionality of Proloquo2Go and Proloquo4Text. Proloquo is now a button-based layout similar to Proloquo2Go but without symbols (only words) and with more predictive text options. Proloquo4Text consists primarily of on-screen typing with the option to save phrases.